# Diffusion-Aided Joint Source Channel Coding For High Realism Wireless Image Transmission

### v2

@@ Abstract
Deep learning-based joint source-channel coding (deep JSCC) has been demonstrated to be an effective approach for wireless image transmission. Nevertheless, most existing work adopts an autoencoder framework to optimize conventional criteria such as Mean Squared Error (MSE) and Structural Similarity Index (SSIM) which do not suffice to maintain the perceptual quality of reconstructed images. Such an issue is more prominent under stringent bandwidth constraints or low signal-to-noise ratio (SNR) conditions. To tackle this challenge, we propose DiffJSCC, a novel framework that leverages the prior knowledge of the pre-trained Statble Diffusion model to produce high-realism images via the conditional diffusion denoising process. Our DiffJSCC first extracts multimodal spatial and textual features from the noisy channel symbols in the generation phase. Then, it produces an initial reconstructed image as an intermediate representation to aid robust feature extraction and a stable training process. In the following diffusion step, DiffJSCC uses the derived multimodal features, together with channel state information such as the signal-to-noise ratio (SNR), as conditions to guide the denoising diffusion process, which converts the initial random noise to the final reconstruction. DiffJSCC employs a novel control module to fine-tune the Stable Diffusion model and adjust it to the multimodal conditions. Extensive experiments on diverse datasets reveal that our method significantly surpasses prior deep JSCC approaches on both perceptual metrics and downstream task performance, showcasing its ability to preserve the semantics of the original transmitted images. Notably, DiffJSCC can achieve highly realistic reconstructions for 768$\times$512 pixel Kodak images with only 3072 symbols ($<$0.008 symbols per pixel) under 1dB SNR channels. The source code will be available at https://github.com/mingyuyng/DiffJSCC.
@@ I Introduction
Robust wireless image and video transmission have become crucial for emerging applications such as augmented reality (AR), virtual reality (VR), and autonomous driving, which commonly offload computation to remote servers. Conventional wireless image transmission utilizes a two-step scheme that first applies compression using codecs like Joint Photographic Experts Group (JPEG) or Better Portable Graphics (BPG), followed by a separate channel coding scheme such as a low-density parity-check code (LDPC) or Polar code for error protection. Nonetheless, these standard compression algorithms are prone to any bit errors and susceptible to dramatic quality degradation (so-called “cliff effect”) when channel coding fails to correct codeword errors entirely. To address this challenge, joint source-channel coding (JSCC) has been proposed to enhance error resilience and maintain image integrity during wireless transmission [1, 2].
Figure 1: Top: End-to-end deep JSCC schemes. Bottom: The proposed JSCC scheme using conditional diffusion models.
Recent advancements in Deep Neural Networks (DNNs) have significantly impacted fields including computer vision [3, 4], natural language processing [5, 6], and wireless communications [7, 8, 9], showcasing their ability to tackle intricate issues including JSCC. Bourtsoulatze et al. [10] have introduced a groundbreaking deep learning-based JSCC (deep JSCC) framework that unifies image compression and error correction through an autoencoder architecture. This holistic end-to-end design, showcased in Figure 1 top, surpasses conventional separate source-channel coding transmission techniques, especially in challenging wireless environments characterized by additive white Gaussian noise (AWGN) and Rayleigh flat fading channels.
@@
Simultaneously, research has developed methods for enhancing deep JSCC’s adaptability to varying channel conditions. Xu et al. developed ADJSCC [20], which employs a Squeeze-Excitation structure [21] to adaptively respond to signal-to-noise ratio (SNR) variations. Yang et al. [22] introduced a policy network designed to modulate transmission rates in accordance with channel status and image contents. Dai et al. [23] optimized rate adjustment using a learned entropy model, and Zhang et al. [24] leveraged a predictive oracle network for informed rate selection.
@@
Although these prior innovations mark significant progress, their focus on reducing image distortion through metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) does not necessarily translate to maintaining perceptual quality, particularly under bandwidth constraints or adverse channel conditions where images may lose texture and structural clarity. Image compression and restoration studies have long faced the issue of blurry reconstructions. Blau et al. [25] highlighted the perception-distortion trade-off, revealing that optimizing rate-distortion criteria alone does not guarantee the preservation of natural image characteristics. To close this gap, integrating adversarial loss into generative models has become prevalent, substantially improving the visual quality of compressed images [26, 27]. With the advent of diffusion models [28] that excel in producing realistic textures in image generation, various diffusion models are now being applied to enhance the quality of reconstructed images even under high compression scenarios, promising a new frontier in image compression [29, 30].
@@
Building on prior research, Yang et al. [13] and Wang et al. [31] incorporated adversarial loss using PatchGAN [32] into deep JSCC frameworks, aiming to improve visual quality. Despite their efforts, the performance was hindered by the difficulties inherent in training effective Generative Adversarial Networks (GANs) from scratch. In a different approach, Ecenaz et al. [33] leveraged the pre-trained StyleGAN-v2 generator [34] to enhance the perceptual quality of transmitted facial images through deep JSCC. However, the reliance on traditional distortion metrics such as mean squared error (MSE) during optimization tempered the gains in practice. Yilmaz et al. [35] recently pioneered the integration of deep JSCC with pixel space pre-trained diffusion models based on the DDNM framework [36], achieving marked improvements in distortion and perceptual quality compared to Ecenaz et al.’s results [33]. Their method separates the range-space from the null-space image components and transmits only the former through deep JSCC while reconstructing the latter with the diffusion model at the receiver. However, since the range-space of the reconstructed image is generated exclusively by deep JSCC, this approach struggles to achieve high-quality image reconstructions when the transmitted range-space components undergo significant distortion, especially in challenging channel conditions.
Recent advancements have optimized the use of pre-trained diffusion models through techniques like ControlNet [37] and T2I-adaptor [38]. These methods customize condition diffusion models on top of the Stable Diffusion model [39] by integrating the Stable Diffusion denoiser with additional lightweight control modules or adapters. These enhancements support various conditions that are not addressed in thte original Stable Diffusion model, including poses, edges, and segmentation masks. Drawing inspiration from these advances, recent work in image restoration [40] and compression [41, 29] has employed similar architectures. These methods capitalize on image content-related features as conditions and apply conditional diffusion generation to reconstruct the final image, demonstrating state-of-the-art performance compared with prior works such as DDNM[36].
Leveraging the latest breakthroughs, we introduce DiffJSCC, a new framework devised to improve the perceptual quality of wirelessly transmitted images through conditional diffusion models as depicted in Figure 1 bottom. Diverging from conventional end-to-end deep JSCC methods, our technique begins by deriving meaningful multimodal features from the noisy output, subsequently utilizing these features to guide a conditional diffusion model based on the pre-trained Stable Diffusion architecture. Our DiffJSCC framework achieves robust feature extraction by utilizing a JSCC decoder for initial image reconstruction, which facilitates the independent training of the encoder and the recovery of strong spatial features. Leveraging the text-to-image synthesis capabilities of Stable Diffusion, we also extract textual features directly from the initial reconstruction using a pre-trained extractor. Additionally, we enhance model performance by integrating channel-specific information, such as SNR, into the conditioning process of diffusion. This multimodal approach allows for effective feature extraction without developing new neural networks for mapping channel outputs to diffusion conditions. Our approach leverages Lin et al.’s control module [40] to adapt the prior knowledge in the pre-trained Stable Diffusion model. The denoising procedure is aligned with the original denoising diffusion probabilistic model (DDPM) [28]. We further refine the process by using the initial image reconstruction as a reference to balance realism with fidelity. Empirical evaluations across diverse image datasets reveal that DiffJSCC surpasses traditional separate source-channel coding schemes and prior deep JSCC benchmarks, showing superior results in human-perceptual quality and practical task efficacy, particularly in scenarios of restricted bandwidth and challenging channel conditions.
@@
Figure 2: The general framework of the proposed DiffJSCC system. In the transmitter segment, the image $x$ is encoded by a JSCC encoder to yield transmission symbols $y$. In the receiver segment, a preliminary image reconstruction $\hat{x}$ is initially generated by a JSCC decoder. Thereafter, $\hat{x}$ is subjected to a robust condition generation to produce multimodal conditions. During the Conditional Latent Diffusion Process, these multimodal attributes are fed to the conditional denoiser to guide the generation procedure.
@@
  • Novel Diffusion-based JSCC Framework: We present a novel JSCC framework that utilizes a pre-trained text-to-image Stable Diffusion model to produce realistic image reconstructions. We develop a custom control module to adapt the Stable Diffusion model for particular datasets unseen during the training. To the best of our knowledge, this is the first instance of integrating the Stable Diffusion model into deep JSCC.
  • Enhanced Multimodal Conditioning: Our framework uniquely extracts multimodal conditions to better guide the Stable Diffusion model by including spatial and textual features and the channel condition information. We demonstrate that these diverse conditions substantially improve performance.
  • Balancing Fidelity and Realness: We propose a denoising guidance process that derives intermediate features with an initial JSCC reconstruction. This approach adeptly balances image fidelity with perceptual realness.
  • Flexible and Versatile Design: DiffJSCC’s model-agnostic design is compatible with various advanced JSCC architectures. Besides, DiffJSCC uses a lightweight transmitter-side model, suiting mobile and IoT applications that offload heavy computations to remote servers.
@@ II Proposed Method
II-A Overview
The architecture of our proposed DiffJSCC framework is illustrated in Figure 2. The input image is denoted as $x\in\mathbb{R}^{C\times H\times W}$, with $C$ indicating the number of color channels and $H$ and $W$ denoting the height and width, respectively. Within the transmitter, a JSCC encoder, $E_{c}$, encodes the image $x$ into a sequence of complex-valued channel input symbols $y\in\mathbb{C}^{K}$. These $K$ symbols are subsequently transmitted through a noisy communication channel while each symbol utilizes the channel once. The rate of the transmission is quantified as the channel utilization per pixel (cpp), defined by the ratio:
$R=\frac{K}{HW}\;\;\text{cpp}.$ (1)
@@ II-A Overview
The wireless channel in our study is modeled as a conditional distribution, with the received symbols being drawn according to $\hat{y}\sim p(y|\varepsilon)$, where $\varepsilon$ encapsulates the channel parameters. We focus on the additive white Gaussian noise (AWGN) channel, which dictates that received symbols $\hat{y}$ are obtained by the following relation:
$\hat{y}=y+n,\;\;n\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{K}).$ (3)
Here, the noise $n\in\mathbb{C}^{K}$ is assumed to have a circularly symmetric complex Gaussian distribution with $\sigma^{2}$ symbolizing the noise’s power level. The channel’s SNR, denoted as $\gamma$, is defined by the ratio $\gamma=\frac{\bar{P}}{\sigma^{2}}$. In our experiments, we standardize the average power $\bar{P}$ to unity for consistency. Aligned with the approach of Xu et al..[20], this study dismisses the slow Rayleigh Fading model under the basis of its effective equivalence to the AWGN channel given perfect channel estimation and equalization. Nevertheless, our proposed methodology is versatile and can be adapted with minimal effort to accommodate alternative channel models. We adhere to the established convention, as presented in preceding studies[20, 22], that the channel SNR is known a priori at both the transmitting and receiving ends. This assumption simplifies the model’s architecture, enabling a single unified model to reliably handle a span of channel SNR conditions.
At the receiver, image reconstruction is structured into two core stages: Robust Condition Generation (Section II-B) and Conditional Latent Diffusion Process (Section II-C). Initially, during the Robust Condition Generation stage, resilient multimodal features are extracted from the noisy received channel symbols to accurately reflect characteristics of the original image and channel state. Subsequently, the Conditional Latent Diffusion Process takes these features as conditioning inputs to a latent diffusion model (LDM). Leveraging ControlNet’s architecture [37], our framework adopts the pre-trained Stable Diffusion model augmented with a control module that fine-tunes the denoising operation for specific conditions, offering high-realism image reconstructions.
II-B Robust Condition Generation
II-B1 Initial JSCC Reconstruction
Rather than directly using the received symbols $\hat{y}$ as conditioning inputs for the conditional latent diffusion model, an intermediate step utilizes a JSCC decoder to obtain an initial image reconstruction. This approach presents three primary benefits: Firstly, it enables separate training of the JSCC encoder and the conditional denoiser, which allows more stable training compared to a joint training approach. Secondly, the initial reconstruction $\hat{x}$ harnesses the pre-trained LDM encoder to capture salient spatial features, thus avoiding the need to develop an encoder from scratch as originally proposed by the ControlNet framework[37]. Lastly, the initial reconstruction $\hat{x}$ facilitates the extraction of supplementary textual descriptors using pre-trained image captioning models such as BLIPv2[42] and the CLIP encoder[43], enriching the input conditions for the diffusion model.
For the initial reconstruction, we employ a JSCC architecture akin to the one described in Yang et al.[22]. The encoder and decoder of the JSCC are trained in unison, where the encoding ($E_{c}$) and decoding ($D_{c}$) processes are expressed as:
$y=E_{c}(x,\gamma),\;\;\hat{x}=D_{c}(\hat{y},\gamma).$ (4)
@@ II-B1 Initial JSCC Reconstruction
$\mathcal{L}_{JSCC}=\mathbb{E}_{x,n}\left[\|x-D_{c}(E_{c}(x,\gamma)+n,\gamma)\|_{2}^{2}\right],$ (5)
where $n$ is the channel noise introduced during transmission. A key aspect of our approach is its model-agnosticism, facilitating seamless integration with more advanced deep JSCC frameworks. While we demonstrate the effectiveness of DiffJSCC using one JSCC structure, the design is inherently adaptable to other more recent architectures such as SwinJSCC[18] or Mamba-JSCC[19].
Figure 3: Network structure of the proposed denoiser $\epsilon_{\theta}$ in the conditional latent diffusion model. The pre-trained Stable Diffusion model is shown in blue and the additional control module is shown in green.
II-B2 Robust Feature Extraction
Leveraging the initial reconstruction $\hat{x}$, our framework extracts multimodal features with the help of pre-trained models, eliminating the need for additional training specific to our approach. The spatial features $f_{spa}$ are extracted by the encoder from the pre-trained Stable Diffusion model, represented as $E_{ldm}$:
$f_{spa}=E_{ldm}(\hat{x}).$ (6)
Concurrently, we use the pre-trained image captioning model BLIPv2 [42] denoted as $E_{blip}$ to generate textual descriptions for $\hat{x}$. Following the initial captioning step, these descriptions are further encoded by the pre-trained CLIP encoder $E_{clip}$[43] to obtain textual features $f_{txt}$:
$f_{txt}=E_{clip}(E_{blip}(\hat{x})).$ (7)
Despite potential quality variations in captions from the initial reconstructions, their inclusion is justified as they encapsulate content-related keywords, proving essential for context preservation.
In addition to multimodal features of $f_{spa}$ and $f_{txt}$, the model integrates relevant channel state information represented solely by the channel SNR $\gamma$ for the AWGN channel considered in this study. Therefore, the set of conditions $\bar{c}$ utilized by our conditional diffusion model is formulated as:
$\bar{c}=\{f_{spa},f_{txt},\gamma\}.$ (8)

### v3

@@ Abstract
Deep learning-based joint source-channel coding (deep JSCC) has been demonstrated to be an effective approach for wireless image transmission. However, many current approaches utilize an autoencoder framework to optimize conventional metrics such as Mean Squared Error (MSE) and Structural Similarity Index (SSIM), which are inadequate for preserving the perceptual quality of reconstructed images. Such an issue is more prominent under stringent bandwidth constraints or low signal-to-noise ratio (SNR) conditions. To tackle this challenge, we propose DiffJSCC, a novel framework that leverages the prior knowledge of the pre-trained Stable Diffusion model to produce high-realism images via the conditional diffusion denoising process. First, our DiffJSCC employs an autoencoder structure similar to prior deep JSCC works to generate an initial image reconstruction from the noisy channel symbols. This preliminary reconstruction serves as an intermediate step where robust multimodal spatial and textual features are extracted. In the following diffusion step, DiffJSCC uses the derived multimodal features, together with channel state information such as the signal-to-noise ratio (SNR) and channel gain. These serve as conditions to guide the diffusion denoising process through a novel control module, which adjusts the output according to the multimodal conditions. To maintain the balance between realism and fidelity, an optional intermediate guidance approach using the initial image reconstruction is implemented. Extensive experiments on diverse datasets reveal that our method significantly surpasses prior deep JSCC approaches on both perceptual metrics and downstream task performance, showcasing its ability to preserve the semantics of the original transmitted images. Notably, DiffJSCC can achieve highly realistic reconstructions for 768$\times$512 pixel Kodak images with only 3072 symbols ($<$0.008 symbols per pixel) under 1dB SNR channels. The source code is available at https://github.com/mingyuyng/DiffJSCC.
@@ I Introduction
Robust wireless image and video transmission have become crucial for emerging applications such as augmented reality (AR), virtual reality (VR), and autonomous driving, which commonly offload computation to remote servers. Conventional wireless image transmission utilizes a two-step scheme that first applies compression using codecs like Joint Photographic Experts Group (JPEG) or Better Portable Graphics (BPG), followed by a separate channel coding scheme such as a low-density parity-check code (LDPC) or Polar code for error protection. Nonetheless, these standard compression algorithms are prone to any bit errors and susceptible to dramatic quality degradation (so-called “cliff effect”) when channel coding fails to correct codeword errors entirely. To address this challenge, joint source-channel coding (JSCC) has been proposed to improve error resilience and maintain image integrity during wireless transmission [1, 2].
Figure 1: Overall framework of the proposed DiffJSCC (e) and comparison with other existing deep JSCC structures (a-d).
Recent advancements in Deep Neural Networks (DNNs) have significantly impacted fields including computer vision [3, 4, 5], natural language processing [6, 7], and wireless communications [8, 9, 10], demonstrating their power in addressing complex problems including JSCC. Bourtsoulatze et al. [11] have introduced a groundbreaking deep learning-based JSCC (deep JSCC) framework that unifies image compression and error correction through an autoencoder architecture. This holistic end-to-end design, shown in Figure 1a, surpasses conventional separate source-channel coding transmission techniques, especially in challenging wireless environments characterized by the Additive White Gaussian Noise (AWGN) channel and the Rayleigh Fading channel.
@@
Simultaneously, research has advanced techniques to enhance deep JSCC’s adaptability to varying channel conditions. Xu et al. developed ADJSCC [21], which employs a Squeeze-Excitation structure [22] to adaptively respond to signal-to-noise ratio (SNR) variations. Yang et al. [23] introduced a policy network designed to modulate transmission rates in accordance with various channel statuses and image contents. Dai et al. [24] optimized rate adjustment using a learned entropy model, and Zhang et al. [25] leveraged a predictive oracle network for informed rate selection.
@@
Although these prior innovations mark significant progress, their focus on minimizing image distortion through metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) does not necessarily translate to maintaining perceptual quality, particularly under bandwidth constraints or adverse channel conditions where images may lose texture and structural clarity. Image compression and restoration studies have long faced the issue of blurry reconstructions. Blau et al. [26] highlighted the perception-distortion trade-off, revealing that optimizing rate-distortion criteria alone does not guarantee the preservation of natural image characteristics. To close this gap, integrating adversarial loss into generative models has become prevalent, substantially improving the visual quality of compressed images [27, 28]. With the advent of diffusion models [29] that excel in producing realistic textures in image generation, various diffusion models are now being applied to enhance the quality of reconstructed images even under high compression scenarios, promising a new frontier in image compression [30, 31].
@@
The application of generative deep learning models has recently been explored in the domain of deep JSCC. Building on prior research, Yang et al. [32] and Wang et al. [33] incorporated adversarial loss using PatchGAN [34] into deep JSCC frameworks (shown in Figure 1.b), aiming to improve visual quality. Despite their efforts, the performance was hindered by the difficulties inherent in training effective Generative Adversarial Networks (GANs) from scratch. In a different approach, Ecenaz et al. [35] leveraged the pre-trained StyleGAN-v2 generator [36] to enhance the perceptual quality of transmitted facial images through deep JSCC (shown in Figure 1.c). However, the reliance on traditional distortion metrics such as mean squared error (MSE) during optimization tempered the gains in practice. Yilmaz et al. [37] recently pioneered the integration of deep JSCC with pre-trained pixel-space diffusion models based on the zero-shot image restoration framework DDNM [38] (as shown in Figure 1.d), achieving notable improvements in distortion and perceptual quality compared to Ecenaz et al.’s results [35]. Their approach involves initially generating an image reconstruction using a conventional JSCC autoencoder. Subsequently, this reconstructed image is utilized to guide the denoising process of the unconditional diffusion model via Range-Null space decomposition. However, since the diffusion process is regulated solely by the intermediate guidance, this method struggles to achieve high-quality image reconstructions, especially when the initial image reconstruction undergoes significant distortion under challenging channel conditions. Additionally, such a zero-shot generation process only supports images as the unimodal condition.
@@
Figure 2: The general framework of the proposed DiffJSCC. In the transmitter, the image $x$ is encoded by a JSCC encoder to yield transmission symbols $y$. In the receiver, a preliminary image reconstruction $\hat{x}$ is generated by a JSCC decoder, which is used to produce multimodal (visual and text) conditions. The final step is the conditional latent diffusion process, which uses these multimodal conditions in the conditional denoiser to guide the image generation procedure.
Recent progress in image restoration [39] and compression [40, 30] has shifted away from solely depending on intermediate guidance for harnessing the prior knowledge in pre-trained diffusion models. Instead, these advancements explore the direct learning of conditional diffusion models through approaches like ControlNet [41] and T2I-adaptor [42], demonstrating state-of-the-art performance and surpassing prior works such as DDNM[38]. These techniques involve fine-tuning the pre-trained diffusion model with additional lightweight control modules or adapters, and considerably reduce the dependence on intermediate guidance. In addition, these fine-tuning methods offer potential support for multimodal conditions to further enhance the control over the image generation process.
Leveraging the latest breakthroughs, we introduce DiffJSCC, a new framework devised to improve the perceptual quality of wirelessly transmitted images through conditional diffusion models as depicted in Figure 1.e. Similar to Yilmaz et al’s method [37], our approach begins with initial image reconstruction via a standard autoencoder framework. In contrast, DiffJSCC focuses on directly learning a conditional diffusion model by fine-tuning the existing pre-trained diffusion model. On one hand, the distorted initial image reconstruction is treated as an input to the conditional denoiser instead of directly guiding the diffusion process, which ensures high-quality image generation even if the initial image reconstruction is highly distorted. On the other hand, this fine-tuning approach allows for multimodal conditions beyond the reconstructed image. Within DiffJSCC, we include text and channel state information (CSI) as additional conditions alongside the visual condition. Both visual and textual conditions are derived via pre-trained neural networks, obviating the need to develop new neural architectures and simplifying the training process. To fuse these multimodal conditions, DiffJSCC utilizes a controller module similar to that in Lin et al. [39]. The benefit of incorporating supplementary textual and CSI conditions has been demonstrated through our experiments. Instead of using pixel-domain diffusion models as in [37], DiffJSCC adopts the advanced latent diffusion model known as Stable Diffusion [43]. The denoising procedure is aligned with the original denoising diffusion probabilistic model (DDPM) [29]. Since DiffJSCC directly learns a conditional diffusion model, intermediate guidance is no longer necessary. However, we demonstrate that this mild intermediate guidance remains beneficial for balancing image realism with fidelity. Empirical evaluations across diverse image datasets reveal that DiffJSCC surpasses traditional separate source-channel coding schemes and prior deep JSCC benchmarks, showing superior results in human-perceptual quality and practical task efficacy, particularly in scenarios of restricted bandwidth and challenging channel conditions.
@@
  • Novel Diffusion-based JSCC Framework: We present a novel JSCC framework that utilizes a pre-trained text-to-image Stable Diffusion model to produce realistic image reconstructions. We develop a custom control module to adapt the Stable Diffusion model for conditional generation for particular datasets. To the best of our knowledge, this is the first instance of integrating the Stable Diffusion model into deep JSCC.
  • Enhanced Multimodal Conditioning: Our framework uniquely extracts multimodal conditions to better guide the Stable Diffusion model by including visual and textual features and the channel information. We demonstrate that these diverse conditions substantially improve the performance.
  • Balancing Fidelity and Realness: We present a denoising guidance process that derives intermediate features with an initial JSCC reconstruction. This approach adeptly balances the image fidelity with perceptual realness.
  • Flexible and Versatile Design: DiffJSCC’s model-agnostic design is compatible with various advanced deep JSCC architectures for initial image reconstruction. Besides, DiffJSCC omits any additional components at the transmitter, suiting mobile and IoT applications that offload heavy computations to remote servers.
@@ II Proposed Method
The general architecture of our proposed DiffJSCC is illustrated in Figure 2, which contains three main components. First, we perform Initial Deep JSCC Reconstruction (Sec, II-A) using traditional autoencoder structures. Then, we perform Multimodal Condition Generation (Sec. II-B) to extract multimodal spatial and textual features from the initial JSCC reconstruction. Lastly, the multimodal features, together with the channel state information, are treated as the conditions and fed to the Conditional Latent Diffusion Process (Sec. II-C) to generate the final refined image reconstruction. More details about the three components are introduced in the following sections.
II-A Initial Deep JSCC Reconstruction
The input image is denoted as $x\in\mathbb{R}^{C\times H\times W}$, with $C$ indicating the number of color channels, and $H$ and $W$ representing the height and width, respectively. Within the transmitter, a JSCC encoder, $E_{c}$, encodes the image $x$ into a sequence of complex-valued channel input symbols $y\in\mathbb{C}^{K}$. These $K$ symbols are subsequently transmitted through a noisy communication channel while each symbol utilizes the channel once. The rate of the transmission is quantified as the channel bandwidth ratio (CFR), which is referred as $\rho$ and is defined as:
$\rho=\frac{K}{C\times H\times W}.$ (1)
@@ II-A Initial Deep JSCC Reconstruction
The wireless channel in our study is modeled as a conditional distribution, with the received symbols being drawn according to $\hat{y}\sim p(y|\varepsilon)$, where $\varepsilon$ encapsulates the channel parameters. In this work, we consider both the AWGN channel and Rayleigh Fading channel, and the channel transfer function can be represented as:
$\hat{y}=hy+n,\;\;n\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{K}).$ (3)
Here, the noise $n\in\mathbb{C}^{K}$ is assumed to have a circularly symmetric complex Gaussian distribution with $\sigma^{2}$ symbolizing the noise’s power level. $h\in\mathbb{C}$ denotes the channel gain. For AWGN channel, $h$ is set to 1. For the Rayleigh Fading channel, the channel gain $h$ is a random variable and it follows $h\sim\mathcal{CN}(0,H)$. The channel’s SNR, denoted as $\gamma$, is defined by the ratio $\gamma=\bar{P}\mathbb{E}[|h|^{2}]/\sigma^{2}$. In our experiments, we standardize the average power $\bar{P}$ to 1 for consistency. Following the prior works [21, 44], both the transmitter and receiver are assumed to have perfect knowledge of the channel gain $h$ and SNR $\gamma$. Thus, at the receiver side, we first perform equalization to the received symbols $\hat{y}$, which is:
$\tilde{y}=\frac{h^{*}}{|h|^{2}}\hat{y},$ (4)
where $h^{*}$ denotes the complex conjugate of $h$.
After that, a JSCC decoder ($D_{c}$) takes the equalized channel symbols $\tilde{y}$ as input and generate the initial reconstruction $\hat{x}\in\mathbb{R}^{C\times H\times W}$. In this paper, we employ a JSCC architecture akin to the one described in Yang et al.[23]. The encoder and decoder could take the CSI information $\gamma$ and $h$ and are trained jointly. The encoding and decoding processes are expressed as:
$y=E_{c}(x,h,\gamma),\;\;\hat{x}=D_{c}(\hat{y},h,\gamma).$ (5)
@@
$\mathcal{L}_{JSCC}=\mathbb{E}_{x,h,n}\left[\|x-\hat{x}\|_{2}^{2}\right],$ (6)
A key aspect of our approach is its model-agnosticism, facilitating seamless integration with more advanced deep JSCC frameworks. While we demonstrate the effectiveness of DiffJSCC using one JSCC structure, the design is inherently adaptable to other more recent architectures such as SwinJSCC[19] or NTSCC[24].
Rather than directly using the received symbols $\hat{y}$ as conditioning inputs for the conditional latent diffusion model, the initial JSCC reconstruction acts as an intermediate step and exhibits three primary benefits: First, it enables separate training of the JSCC encoder and the conditional denoiser, which allows more stable training compared to a joint training approach. Second, the initial reconstruction $\hat{x}$ harnesses the pre-trained LDM encoder to capture salient spatial features, thus avoiding the need to develop an encoder from scratch as originally proposed by the ControlNet framework[41]. Last, the initial reconstruction $\hat{x}$ facilitates the extraction of supplementary textual descriptors using pre-trained image captioning models such as BLIPv2[45] and the CLIP encoder[46], enriching the input conditions for the diffusion model.
II-B Multimodal Condition Generation
Leveraging the initial reconstruction $\hat{x}$, our framework extracts multimodal features with the help of pre-trained neural networks, eliminating the need for additional training specific to our approach. The spatial features $f_{v}$ with a size of $84\times H/8\times W/$ 8 are extracted by the image encoder $E_{ldm}$ from the pre-trained Stable Diffusion model:
$f_{v}=E_{ldm}(\hat{x}).$ (7)
In parallel, given that the Stable Diffusion model is inherently a text-to-image generative model, it is natural to leverage textual information as a supplementary condition. Noticing the presence of an initial image reconstruction, we propose to utilize the pre-trained image captioning model BLIPv2 [45] denoted as $E_{blip}$ to generate textual descriptions for $\hat{x}$. Following the initial captioning step, these descriptions are further encoded by the pre-trained CLIP encoder $E_{clip}$[46] to obtain textual features $f_{t}\in\mathbb{R}^{77\times 768}$:
$f_{t}=E_{clip}(E_{blip}(\hat{x})).$ (8)
An alternative approach to leveraging textual data is by deploying image captioning models at the transmitter end and sending the generated captions along with the image. Nevertheless, since image captioning models often involve significant computational requirements, this method could substantially increase the transmitter’s processing load, making it impractical for mobile systems. It also incurs the overhead of sending text captions through the noisy channel. Within our method, despite the potential quality degradation in captions due to the distortion of initial image reconstructions, the inclusion of $f_{t}$ generated from $\hat{x}$ has been shown to be beneficial as it encapsulates content-related keywords.
The CSI information, described by the channel SNR $\gamma$ and the channel gain $h$, plays a crucial role alongside the multimodal features $f_{v}$ and $f_{t}$ as it provides insight into the extent of distortion that initial image reconstruction experiences. After integrating the CSI, the set of conditions $\bar{c}$ utilized by our conditional diffusion model is formulated as:
$\bar{c}=\{f_{v},f_{t},h,\gamma\}.$ (9)